OpenStudy (anonymous):

the radius of a circle is 3.5ft , the height is 14ft what is the volume and surface area?

OpenStudy (wolf1728):

Basically, you are asking for the area of a cylinder right?

OpenStudy (kc_kennylau):

Cylinder or Cone?

OpenStudy (anonymous):

yes surface area and volume

OpenStudy (anonymous):

cylinder

OpenStudy (kc_kennylau):

V=pi*r^2*h A=2*pi*r(r+h)

OpenStudy (wolf1728):

Volume = 2 * PI * r * height Volume = 2 * 3.14159 * 3.5 * 14 Volume = 307.88

OpenStudy (anonymous):

cool thats what i got but im not sure about the surface area

OpenStudy (wolf1728):

Should be r²

OpenStudy (anonymous):

quickly wolf its 3 am over here

OpenStudy (wolf1728):

Volume = 3.14159 * 3.5² * 14 = 538.78

OpenStudy (anonymous):

thanks man

OpenStudy (wolf1728):

Need the area right?

OpenStudy (anonymous):

yeah

OpenStudy (wolf1728):

TOTAL Surface Area = 384.85 I used calculator here: http://1728.org/diam.htm

OpenStudy (anonymous):

thanks man your the best

OpenStudy (wolf1728):

u r welcome :-)